OpenStudy (anonymous):

let f(x) = 1 - 4x^2. transform f into vertex form. find the y intercept x intercept and the vertex of the function

OpenStudy (anonymous):

Vertx: (0,1) x-int : y-int: 0=1-4x^2 y= 1-4(0)^2 -1 = -4x^2 y = 1 \[-\frac{1}{-4}=x^2\] \[x=\pm \sqrt{\frac{1}{4}}\]\[x=\pm\frac{1}{2}\]